OpenStudy (anonymous):

It's given that V(t)=Vo+at Vo= 25.00 m/s V=17. m/s after .400s and 3.00 m/s after 1.100s What is the acceleration? I know it is going to be m/s^2, but that's what's confusing me because I've worked it out a number of ways and never have had to square it. If someone could show me how to work it that would be awesome.

OpenStudy (anonymous):

(1) 17 = 25+ a x 0.4 (2) 3 = 25 + a x 1.100

OpenStudy (anonymous):

solve both these equations fore a and then equate

OpenStudy (anonymous):

wow thank you so much

OpenStudy (anonymous):

(1)* .......a= -8/0.4 (2)*............a=-22/1.100

OpenStudy (anonymous):

its a deacceleration thats why a is negative notice how the velocity is slowing down as the time increases, must be applying the break

OpenStudy (anonymous):

a=-20 m/s^2

OpenStudy (anonymous):

But why is acceleration squared?

OpenStudy (anonymous):

That's what I did to get to that point, but it just confused me because I didn't square anything so acceleration shouldn't have been m/s^2 algebraically.

OpenStudy (anonymous):

units acceleration is the rate of change of velocity per unit time